# Characterization of Silicon Photomultiplier Photon Detection Efficiency at Liquid Nitrogen Temperature

### v1

@@ Abstract
The detection of individual photons at cryogenic temperatures is of interest to many experiments searching for physics beyond the Standard Model. Silicon photomultipliers are often deployed in liquid argon or liquid xenon to detect scintillation light by either directly detecting the vacuum ultra-violet scintillation or by detecting light from fluorescent compounds that are used to shift the wavelength. Here we present results from an experimental setup that measures the photon detection efficiencies of silicon photomultipliers, sensitive to the visible spectrum, at liquid nitrogen temperature, 77 K. Results from a KETEK PM3325-WB-D0 and a Hamamatsu S13360-3050CS silicon photomultiplier exhibit a decrease in photon detection efficiency greater than 20% at liquid nitrogen temperature relative to room temperature for $562.5$ nm light.
@@ 1 Introduction
A silicon photomultiplier (SiPM) is a solid-state detector with single-photon resolution and can be used to determine the arrival time of individual photons with nanosecond precision. This makes SiPMs ideal candidates for low-light detection applications [1]. A SiPM contains an array of Single Photon Avalanche Diodes (SPADs), each containing a quenching RC circuit, and is operated a few volts above the breakdown voltage. An incident photon can be absorbed within a SPAD and trigger a self-sustaining avalanche of charge carriers with some probability. Several rare-event search experiments have operated SiPMs at cryogenic temperatures [2, 3], or are planning on operating at cryogenic temperatures to detect the scintillation light from liquid argon and liquid xenon [4, 5, 6].
@@
The work here is motivated by LEGEND — an experiment searching for neutrinoless double-beta ($0\nu\beta\beta$) decay in 76Ge — because it uses SiPMs for the readout of its liquid argon (LAr) scintillation detector [7, 8]. The detection of argon scintillation in coincidence with a germanium detector energy deposition is used to tag and remove events that do not resemble $0\nu\beta\beta$ decay, which should be contained entirely within the germanium. The liquid argon scintillation, peaked at 128 nm in the vacuum ultraviolet (VUV), is wavelength-shifted and guided to several arrays of SiPMs by a shroud of tetraphenyl butadiene (TPB) coated, wavelength-shifting optical fibers [9]. This method of collecting the VUV scintillation offers a large area of exposure to scintillation as opposed to using VUV-sensitive SiPMs; in addition, the wavelength-shifting also better matches the spectral response of commercially available visible-light sensitive SiPMs. However, the total collection efficiency of the liquid argon veto — without accounting for shadowing from the LEGEND geometry — is expected to be $\sim 0.1\%$ of the total number of VUV photons produced [10].
@@
Characterizing the photon detection efficiency (PDE) of SiPMs at cryogenic temperatures is important for understanding the LAr scintillation detector response of LEGEND-200 — the current phase of LEGEND that is actively taking data. The PDE is the probability of a SiPM to detect a single incident photon and convert it into a useful signal. Knowledge of the PDE is critical for modeling and understanding the backgrounds observed in the experiment. These cryogenic SiPM characterizations will also inform the design of the future phase of LEGEND with 1000 kg of active isotope which aims to reach a discovery sensitivity for $0\nu\beta\beta$ decay beyond $10^{28}$ years. Knowledge of the SiPMs’ PDE at liquid argon temperature (87 K) from experimental characterizations is important for optimizing LEGEND-1000’s LAr light readout, and for accurately projecting the background [8]. In this paper, we report a characterization of the cryogenic PDEs of a KETEK PM3325-WB-D0 SiPM and a Hamamatsu S13360-3050CS SiPM [^1] LEGEND currently uses KETEK PM33100T SiPMs; the KETEK PM3325-WB-D0 was one of the latest available SiPM models from KETEK before the acquisition of SiPM assets by BROADCOM®. The Hamamatsu SiPM, in contrast, has a higher PDE at the peak emission wavelength of the wavelength-shifting fibers currently used in LEGEND. near liquid argon temperature to determine potential effects on the efficiency of the LEGEND liquid argon scintillation readout system. We find that the cryogenic PDEs at $562.5$ nm for these two devices exhibit a greater than 20% decrease from their room temperature PDE values across a broad range of overvoltages.
@@ 2 Experimental Method
A SiPM characterization test stand has been built at the University of Washington to measure the PDE and other properties, such as the dark count rate and breakdown voltage, at both room temperature and at cryogenic temperature. The PDE ($\varepsilon_{PDE}$) of a SiPM can be broken down into three components [1],
$\varepsilon_{PDE}(V_{OV},\lambda,T)=QE(\lambda,T)\times P_{t}(V_{OV},\lambda,T)\times FF_{\mathrm{eff.}}(V_{OV},\lambda)\,.$ (2.1)
@@
The setup consists of a ThorLabs IS200-4 integrating sphere suspended above a liquid nitrogen bath inside of a dewar contained inside a dark box as shown in Fig. 1. At each port of the integrating sphere, with perpendicular lines of sight, are the SiPM under test, a calibrated BROADCOM AFBR-S4K33C0135L SiPM (the reference SiPM), and a pulsed light emitting diode (LED) light source peaked at 562.5 nm with a full width at half maximum of 25 nm. One NanOpticsTM(030UC-00S) clear, double-cladded optical fiber connects the LED light source —located outside of the dewar — to the integrating sphere. The BROADCOM SiPM is used as a reference for the number of photons in the integrating sphere because its room temperature PDE is known absolutely. For room temperature PDE measurements, the reference diode is mounted directly to the integrating sphere; for cryogenic PDE measurements, the reference SiPM is placed at the lid of the dewar (thereby keeping it at room temperature) and is optically coupled to the integrating sphere by a bundle of the clear fibers. The SiPM signals are read out when triggered by an Agilent 33521 pulse generator that pulses the LED. The SiPM signals are pre-processed first with Texas Instruments LMH6629 transimpedance amplifiers near the integrating sphere and then with pre-amps consisting of two Analog Devices AD8014 amplifiers in series outside the dewar.
@@
The absolute PDE ($\varepsilon_{PDE}^{\mathrm{abs.}}$) at liquid nitrogen temperature is calculated by taking the ratio of the number of detected photons at the SiPM ($N_{SiPM}$) and the incoming flux of photons at the SiPM port ($\Phi_{SiPM}$) multiplied by the SiPM area ($A_{SiPM}$),
$\varepsilon_{PDE}^{\mathrm{abs.}}(V_{OV},\lambda,77~{}\mathrm{K})=\frac{N_{SiPM}(V_{OV},\lambda,77~{}\mathrm{K})}{\Phi_{SiPM}A_{SiPM}}.$ (2.2)
@@
The incident flux on the test SiPM port is assumed to be equal to the incident flux on the reference SiPM port. During liquid nitrogen PDE measurements, the reference SiPM is placed at the lid of the dewar and is optically coupled to the integrating sphere by a bundle of the clear, double-cladded optical fibers, in order to preserve the known response ($\varepsilon_{PDE}^{\mathrm{ref.}}$) of the reference SiPM. Because the LED’s emission spectrum, $\rho(\lambda)$, spans a range of wavelengths, the reference SiPM’s PDE should be a weighted average of the PDE across the LED’s emission spectrum,
@@
$\Phi_{SiPM}=\frac{N_{\mathrm{ref.}}}{A_{\mathrm{ref.}}\varepsilon_{PDE}^{\mathrm{ref.}}f},$ (2.4)
where $N_{\mathrm{ref.}}$ is the number of photons detected by the reference SiPM during the liquid nitrogen temperature measurement, $A_{\mathrm{ref.}}$ is the area of the reference SiPM, and $f$ is the flux ratio that corrects for the transmission coefficient to the reference SiPM (including, but not limited to, optical fiber transmission and coupling to the fibers) and any increased reflectivity in the sphere due to the fibers. The equality in Eqn. 2.4 assumes that light is distributed equally to both integrating sphere ports, that the transmission coefficient does not vary with temperature, and that the SiPMs are aligned perfectly normal to the port. The flux ratio can be measured by first comparing the number of photons measured by the reference SiPM at room temperature in this relative PDE configuration to the number of photons measured at room temperature when the reference SiPM is placed directly on the sphere. This is then normalized by the number of photons detected by the SiPM under test in order to correct for any fluctuations in the LED’s brightness or increased reflectivity from the fibers,
$f=\frac{N_{\mathrm{ref.}}^{\mathrm{rel.}}}{N_{\mathrm{ref.}}^{\mathrm{abs.}}}\frac{N_{\mathrm{SiPM}}^{\mathrm{abs.}}}{N_{\mathrm{SiPM}}^{\mathrm{rel.}}}.$ (2.5)
Here $N_{\mathrm{ref.}}^{\mathrm{abs.}}$ and $N_{\mathrm{SiPM}}^{\mathrm{abs.}}$ are measured when both SiPMs are mounted directly on the integrating sphere, while $N_{\mathrm{ref.}}^{\mathrm{rel.}}$ and $N_{\mathrm{SiPM}}^{\mathrm{rel.}}$ are measured with the test SiPM on the integrating sphere and the reference SiPM at the lid. These four quantities are measured at room temperature.
@@ 3 Data Acquisition and Processing
The signal output of the SiPMs were recorded by a CAEN DT5730 digitizer with a 500 MS/s sampling rate, 14-bit resolution, and dynamic range of 2 Vpp. The data were acquired using CAEN CoMPASS software [11]. The LED was pulsed with a 16 ns pulse width and a 1 kHz frequency from the function generator. Waveforms of length 10 $\mu$s, with the trigger at $1.5\mu$s, were acquired to ensure the end of the waveform window would mimic dark conditions so that a dark count rate could be measured. Three sets of waveform data were taken with 90 s acquisition time at each overvoltage in order to improve statistics. This process of collecting waveform data at each overvoltage was done with the absolute PDE configuration at room temperature, and with the relative PDE configuration at room and liquid nitrogen temperatures. Further dark data were taken to enable a full characterization of the two SiPMs, see appendix A.
To ensure that the PDE measurements were not biased by correlated noise such as cross-talk, the pedestal method of measuring the PDE was used as described in [12]. To summarize, waveforms were acquired synchronously by triggering on the pulser so that baseline events, where the SiPM did not detect photons, were recorded. If the number of detected photons is Poisson distributed, then the probability of not detecting any photoelectrons, $P(N_{\mathrm{detected}}=0)$, gives an unbiased estimate of the Poissonian mean ($\overline{N_{\gamma}}$) because the baselines are unbiased by any correlated noise:
@@
To get the number of baseline events $N_{0}$ the SiPM signals are first integrated in a 1 $\mu$s window around the LED trigger time as well as the last 1 $\mu$s of the waveform to get the dark count rate. Examples of the synchronously triggered waveforms are shown in Fig. 2a. When the histograms of these integrated signals are made, the peaks that correspond to the integral of the baseline events are called the pedestal peaks, as can be seen in Fig. 2b. The light level of the pulsed LED was set low enough so that an appreciable number of baseline events were acquired by making sure the pedestal peak was resolvable by eye when using the CoMPASS online energy spectrum. A Gaussian is then fit to the pedestal peaks in the histograms to determine the pedestal peak locations $\mu$ and standard deviations $\sigma$. The number of baseline events $N_{0}$ in light and dark conditions are then determined by the number of counts below a $\mu+3\sigma$ threshold. This routine is necessary because differences in the electronics noise between room and liquid nitrogen temperatures preclude the use of a single threshold. To get the mean number of photons detected by either the test SiPM or the reference SiPM ($N_{\gamma}$) requires subtraction of the dark count rate,

### v2

@@ Abstract
The detection of individual photons at cryogenic temperatures is of interest to many experiments searching for physics beyond the Standard Model. Silicon photomultipliers (SiPMs) are often deployed in liquid argon or liquid xenon to detect scintillation light either directly or after it has been wavelength-shifted. Maximizing the photon detection efficiency (PDE) of the SiPMs used in these experiments optimizes the sensitivity to new physics; however, the PDEs of commercial SiPMs, although well known at room temperature, are not well characterized at the cryogenic temperatures at which many experiments operate them. Here we present results from an experimental setup that measures the photon detection efficiencies of silicon photomultipliers at liquid nitrogen temperature, 77 K. Results from a KETEK PM3325-WB-D0 and a Hamamatsu S13360-3050CS silicon photomultiplier — of R&D interest to the LEGEND experiment — exhibit a decrease in photon detection efficiency greater than 20% at liquid nitrogen temperature relative to room temperature for $562$ nm light.
@@ 1 Introduction
Silicon photomultipliers (SiPMs) are solid-state detectors with single-photon resolution and can be used to determine the arrival time of individual photons with sub-nanosecond precision due to their good single photon time resolutions and good photon detection efficiencies [1, 2, 3, 4, 5]. This makes SiPMs ideal candidates for low-light detection applications [6, 7, 8, 1, 9, 5]. A SiPM contains an array of Single Photon Avalanche Diodes (SPADs), each containing a quenching resistor, and is operated a few volts above the breakdown voltage. An incident photon can be absorbed within a SPAD and trigger a self-sustaining avalanche of charge carriers with some probability. Several rare-event search experiments have operated SiPMs at cryogenic temperatures in liquid argon [10, 11, 12, 13, 14] and liquid xenon [15], or are planning on operating with SiPMs optimized to detect the scintillation light from liquid argon [16, 17, 18] and liquid xenon [19, 20, 21].
@@
The work here is motivated by LEGEND — an experiment searching for neutrinoless double-beta ($0\nu\beta\beta$) decay in 76Ge — because it uses SiPMs for the readout of its liquid argon (LAr) scintillation detector [22, 23]. The detection of argon scintillation in coincidence with a germanium detector energy deposition is used to tag and remove events that do not resemble $0\nu\beta\beta$ decay, which should be contained entirely within the germanium. Silicon photomultiplers were chosen as the main light readout technology for LEGEND-200 over photomultiplier tubes due to their low radioactive background, low operating voltage, and comparably high quantum efficiency [24, 11, 25]. The liquid argon scintillation, peaked at 128 nm in the vacuum ultraviolet (VUV), is wavelength-shifted and guided to several arrays of SiPMs by a shroud of tetraphenyl butadiene (TPB) coated, wavelength-shifting (WLS) optical fibers [11]. There are two wavelength conversions involved in this collection process: the TPB first shifts the VUV scintillation light to approximately 430 nm [26], and then the WLS optical fibers shift the light to green wavelengths, approximately 460 nm to 580 nm. These green light wavelengths (460 nm to 580 nm) were chosen because they are well matched to the spectral response of the KETEK PM33100T SiPMs used for light readout in LEGEND-200 [27], as well as other models of SiPMs such as the RGB-HD from Fondazione Bruno Kessler, the S13360 from Hamamatsu, or any other commercially available visible light sensitive SiPMs. This method of collecting the VUV scintillation offers a large area of exposure to scintillation as opposed to using VUV-sensitive SiPMs; in addition, the wavelength-shifting also better matches the spectral response of commercially available visible-light sensitive SiPMs. However, the total collection efficiency of the liquid argon detector — without accounting for shadowing from the LEGEND geometry — is expected to be $\sim 0.1\%$ of the total number of VUV photons produced [28].
@@
Characterizing the photon detection efficiency (PDE) of SiPMs at cryogenic temperatures is important for modeling the LAr scintillation detector response of LEGEND-200 — the current phase of LEGEND that is actively taking data. The PDE is the probability of a SiPM detecting a single incident photon and converting it into a useful signal. The overall light collection efficiency of the LAr detector is the product of individual collection, absorption, transmission, emission, and detection probabilities in the chain of steps a VUV photon takes from scintillation to detection in a SiPM; the PDE is the final link in this concatenation. Cryogenic SiPM PDE measurements will also inform the design of the future phase of LEGEND with 1000 kg of active isotope which aims to reach a discovery sensitivity for $0\nu\beta\beta$ decay beyond $10^{28}$ years. Knowledge of the SiPMs’ PDE at liquid argon temperature (87 K) from experimental characterizations is important for optimizing LEGEND-1000’s LAr light readout, and for accurately projecting the background [23].
Although most measurements of the cryogenic PDE of SiPMs tend to focus on SiPMs that are optimized for cryogenic operation [29, 30, 31], a number of measurements of the cryogenic PDE of non-cryogenically-optimized SiPMs have been made and are available [32, 33]. Several of these measurements have shown that the PDE consistently drops by 20% to 30% at cryogenic temperatures compared to room temperature values across a range of wavelengths [32, 29]. The physical mechanism responsible for this decrease is not well understood, but is hypothesized to originate from a decrease in either the quantum efficiency, the avalanche triggering probability, or a combination of the two [32, 33]. A drop in quantum efficiency can be explained by an increase in absorption length, and a drop in avalanche triggering probability can be explained by carrier freeze-out at lower temperatures. The characterization of SiPMs at cryogenic temperatures is important for understanding these effects, but is quite challenging as it requires a dedicated setup that holds only the SiPM at cryogenic temperature while ensuring that other optical properties of the setup do not change. The design of our apparatus cools and maintains the SiPM, at ambient pressure, at 77 K by keeping the SiPM in thermal contact with a simple liquid nitrogen reservoir in a dewar. Other experimental setups rely on vacuum systems to guarantee optical stability, while our approach uses liquid nitrogen boil off to similar effect, as systematic studies in section 4 show. The liquid nitrogen bath only allows us to probe one temperature — unlike other experiments that use closed-cycle cryostats or heat-exchangers — close to LEGEND’s 87 K operating temperature where the results are the most critical to understanding the impact of a changing PDE on the overall light collection efficiency of LEGEND’s liquid argon detector.
In this paper, we describe the design of a test stand capable of measuring the PDE at cryogenic temperatures. We report a characterization of the cryogenic PDEs of two commercially available SiPMs — a KETEK PM3325-WB-D0 SiPM and a Hamamatsu S13360-3050CS SiPM — near liquid argon temperature to determine potential effects on the efficiency of the LEGEND liquid argon scintillation readout system. LEGEND currently uses KETEK PM33100T SiPMs; the KETEK PM3325-WB-D0 was one of the latest available SiPM models from KETEK before the acquisition of SiPM assets by BROADCOM®. The Hamamatsu SiPM, in contrast, has a higher PDE at the peak emission wavelength of the wavelength-shifting fibers currently used in LEGEND and is of R&D interest for LEGEND-1000. We find that the cryogenic PDEs at $562$ nm for these two devices exhibit a greater than 20% decrease from their room temperature PDE values across a broad range of overvoltages.
@@ 2 Experimental Method
A SiPM characterization test stand has been built at the University of Washington to measure the PDE and other properties, such as the dark count rate and breakdown voltage, at both room temperature and cryogenic temperature. The $\mathrm{PDE}$ of a SiPM can be broken down into three components [6],
$\mathrm{PDE}(V_{OV},\lambda,T)=QE(\lambda,T)\times P_{t}(V_{OV},\lambda,T)\times FF_{\mathrm{eff.}}(V_{OV},\lambda)\,.$ (2.1)
@@
The setup consists of a ThorLabs IS200-4 integrating sphere suspended above a liquid nitrogen bath inside of a dewar contained inside a dark box as shown in Fig. 1. At each port of the integrating sphere, with perpendicular lines of sight, are the SiPM under test, a BROADCOM AFBR-S4K33C0135L SiPM (the reference SiPM), and a pulsed light emitting diode (LED) light source peaked at 562 nm with a full width at half maximum of 11 nm [34]. This wavelength was chosen because it is within LEGEND’s WLS fibers’ emission spectrum, approximately 460 nm to 580 nm [27], and was readily at hand. One NanOpticsTM(030UC-00S) clear, double-cladded optical fiber connects the LED light source —located outside of the dewar — to the integrating sphere. The SiPM under test sits on top of a copper cold finger, which allows it to be coupled thermally to a liquid nitrogen bath. The cold finger also ensures that the integrating sphere and the surface of the SiPM under test are not submerged in liquid nitrogen, which would alter the transmissivity and prohibit a clear measurement of the intrinsic PDE; however, future measurements should be done to measure the transmissivity in liquid argon to determine its impact on the overall liquid argon detector efficiency.
The uncalibrated BROADCOM SiPM is used as a reference for the relative light level in the integrating sphere because an estimate of its room temperature PDE is available from the manufacturer and is constant across measurements. For room temperature PDE measurements, the reference diode is mounted directly to the integrating sphere; for cryogenic PDE measurements, the reference SiPM is placed at the lid of the dewar, thermally isolated from the LN, and optically coupled to the integrating sphere by a bundle of clear fibers. The SiPM signals are read out when triggered by an Agilent 33521 pulse generator that pulses the LED. The SiPM signals are pre-processed first with Texas Instruments LMH6629 transimpedance amplifiers near the integrating sphere and then with pre-amps consisting of two Analog Devices AD8014 amplifiers in series outside the dewar, based on the amplifier found in [35]. The Texas Instruments LMH6629 amplifier near the integrating sphere increases the signal-to-noise ratio so that single-photon resolution can be achieved during room temperature measurements.
@@
The PDE at liquid nitrogen temperature is calculated by taking the ratio of the number of detected photons at the SiPM ($N_{SiPM}$) and the incoming flux of photons at the SiPM port ($\Phi_{SiPM}$) multiplied by the SiPM area ($A_{SiPM}$). Calculation of $N_{SiPM}$ is done using Eqn. 3.3 and is detailed in section 3. We are using the reference SiPM to calculate $\Phi_{SiPM}$, and because the reference SiPM has not been absolutely calibrated (we are taking its value from the manufacturer’s datasheet), the PDE we measure is relative to the reference and is given by
$\mathrm{PDE}^{\mathrm{rel.}}(V_{OV},\lambda,77~{}\mathrm{K})=\frac{N_{SiPM}(V_{OV},\lambda,77~{}\mathrm{K})}{\Phi_{SiPM}A_{SiPM}}.$ (2.2)
@@
The incident flux on the test SiPM port is assumed to be equal to the incident flux on the reference SiPM port. During liquid nitrogen PDE measurements, the reference SiPM is placed at the lid of the dewar and is optically coupled to the integrating sphere by a bundle of the clear, double-cladded optical fibers, in order to preserve the known response ($\mathrm{PDE}^{\mathrm{ref.}}$) of the reference SiPM. Because the LED’s emission spectrum, $\rho(\lambda)$, spans a range of wavelengths, the reference SiPM’s PDE should be a weighted average of the PDE across the LED’s emission spectrum,
@@
$\Phi_{SiPM}=\frac{N_{\mathrm{ref.}}}{A_{\mathrm{ref.}}\mathrm{PDE}^{\mathrm{ref.}}f},$ (2.4)
where $N_{\mathrm{ref.}}$ is the number of photons detected by the reference SiPM during the liquid nitrogen temperature measurement, $A_{\mathrm{ref.}}$ is the area of the reference SiPM, and $f$ is the flux ratio that corrects for the transmission coefficient to the reference SiPM (including, but not limited to, optical fiber transmission and coupling to the fibers, which are measured in Section. 4) and any increased reflectivity in the sphere due to the fibers. The equality in Eq. 2.4 assumes that light is distributed equally to both integrating sphere ports, that the transmission coefficient does not vary with temperature, and that the SiPMs are aligned perfectly normal to the port. These assumptions are verified in Section. 4. The flux ratio can be measured by first comparing the number of photons measured by the reference SiPM at room temperature in this lid PDE configuration to the number of photons measured at room temperature when the reference SiPM is placed directly on the sphere. This is then normalized by the number of photons detected by the SiPM under test in order to correct for any fluctuations in the LED’s brightness or increased reflectivity from the fibers,
$f=\frac{N_{\mathrm{ref.}}^{\mathrm{lid}}}{N_{\mathrm{ref.}}^{\mathrm{dir.}}}\frac{N_{\mathrm{SiPM}}^{\mathrm{dir.}}}{N_{\mathrm{SiPM}}^{\mathrm{lid}}}.$ (2.5)
Here $N_{\mathrm{ref.}}^{\mathrm{dir.}}$ and $N_{\mathrm{SiPM}}^{\mathrm{dir.}}$ are measured when both SiPMs are mounted directly on the integrating sphere, while $N_{\mathrm{ref.}}^{\mathrm{lid}}$ and $N_{\mathrm{SiPM}}^{\mathrm{lid}}$ are measured with the test SiPM on the integrating sphere and the reference SiPM at the lid. These four quantities are measured at room temperature and the value of $f$ is of order $\sim 1$.
@@ 3 Data Acquisition and Processing
The signal output of the SiPMs were recorded by a CAEN DT5730 digitizer with a 500 MS/s sampling rate, 14-bit resolution, and dynamic range of 2 Vpp. The data were acquired using CAEN CoMPASS software [36]. The LED was pulsed with a 16 ns pulse width and a 1 kHz frequency from the function generator. Waveforms of length 10 $\mu$s, with the trigger at $1.5\mu$s, were acquired. These waveforms were long enough to ensure that the last microsecond of data — which is several microseconds after the LED has turned off and the SiPM has recharged — is taken in dark conditions so that a dark count rate could be measured. Three sets of waveform data were taken with 90 s acquisition time at each overvoltage in order to improve statistics. This process of collecting waveform data at each overvoltage was done with the direct PDE configuration at room temperature, and with the lid PDE configuration at room and liquid nitrogen temperatures. Further dark data were taken to enable a full characterization of the two SiPMs, see appendix A.
To ensure that the PDE measurements were not biased by correlated noise such as cross-talk, the pedestal method of measuring the PDE was used as described in [37]. To summarize, waveforms were acquired synchronously by triggering on the pulser so that baseline events, where the SiPM did not detect photons, were recorded. If the number of detected photons is Poisson distributed, then the probability of not detecting any photoelectrons, $P(N_{\mathrm{detected}}=0)$, gives an unbiased estimate of the Poissonian mean ($\overline{N_{\gamma}}$) because the baselines are unbiased by any correlated noise [37]:
@@
To get the number of baseline events $N_{0}$ the SiPM signals are first integrated in a 1 $\mu$s window around the LED trigger time as well as the last 1 $\mu$s of the waveform to get the dark count rate. Examples of the synchronously triggered waveforms are shown in Fig. 2a. When the histograms of these integrated signals are made, the peaks that correspond to the integral of the baseline events are called the pedestal peaks, as can be seen in Fig. 2b. The light level of the pulsed LED was set low enough so that an appreciable number of baseline events were acquired by making sure the pedestal peak was resolvable by eye when using the CoMPASS online energy spectrum. This corresponds to roughly $5.73\pm 0.08$ photons per pulse incident on the SiPM face, as measured by the reference diode in the direct PDE configuration of the test stand. A Gaussian is then fit to the pedestal peaks in the histograms to determine the pedestal peak locations $\mu$ and standard deviations $\sigma$. The number of baseline events $N_{0}$ in light and dark conditions are then determined by the number of counts below a $\mu+3\sigma$ threshold. This routine is necessary because differences in the electronics noise between room and liquid nitrogen temperatures preclude the use of a single threshold. To get the mean number of photons detected by either the test SiPM or the reference SiPM ($N_{\gamma}$) requires subtraction of the dark count rate,